A survey of 62 students found that 33% were in favor of raising tution to build a new football stadium. The standard deviation of the sample proprortion is 8.9%. How large a sample (to the nearest person) would be required to reduce this standard deviation to 6.5%? Answer:
Answer #1

Answer:

Given,

Standard deviation = 8.9

n = 62

To determine the sample size

Consider,

variance = p*q / n

8.9^2 = p*q / 62

79.21 = p*q / 62

p*q = 79.21*62

= 4911.02

pq = 4911.02

Now consider standard deviation = 6.5

6.5^2 = p*q / n

42.25 = 4911.02 / n

n = 4911.02 / 42.25

n = 116.2371598

n = 116

Sample size = 116